You don’t have to explain
Just fell the blank
An economy consists of three workers: Lary Moe and Curly. Each works 10 hours a day and can produce and washing hour Larry can either mow 1 lawn or wash 1 car, Moe can either mow 1 lawn or wash 2 cars and Curly can all three spend all their time washing cars, they wil all three spend all their tme mowing lawns, they will all three spend haif their Sme on each acivity, they will wash Larry spends half his time on each activity, whle Moe only washes cars and they will wash lawns which point depicts QUESTION 3 2 points Save Answer In the Larry, Moe and Curly scenario, label the following allocations of time as either efficient (E) o inefficient 0 All three spend all their time mowing lawns All three spend all their time washing cars All three spend half their time on each activity. Larry spends half his time on each activity, while Moe only washes cars and Curly only mows lawns
Answer #1

1).

There are three workers, “Larry”, “Moe” and “Curly” each works 10hours a day. In an hour Larry can produce either mow 1 lawn or wash 1 car; Moe can produce either mow 1 lawn or wash 2 cars and Curly can produce either mow 2 lawns or wash 1 cars.

If all three spend all their time washing cars, they will wash “1*10+2*10+1*10 = 40 cars” and mow “0 lawns”. If all three spend all their time mowing lawns, they will wash “0 cars” and mow “1*10+1*10+2*10 = 40 lawns”.

If all three spend half their time on each activity, they will wash “1*10/2+2*10/2+1*10/2 = 20 cars” and mow “1*10/2+1*10/2+2*10/2 = 20 lawns”.

If Larry spends half his time on each activity, while Moe only washes cars and Curly only mows lawns they will wash “1*10/2+2*10 = 25” cars and mows “1*10/2+2*10=25” lawns.

2).

Consider the PPF. All three spend all their time moving lawns, => it shows the point “E”.

All three spend all their time washing cars, => it shows the point “A”.

All three spend half their time on each activity, => it shows the point “F”. Larry spends half his time on each activity, while Moe only washes cars and Curly only mows lawns, => it shows the point “C”.

3).

All three spend all their time moving lawns, => it is on the PPF, => it’s an efficient point, => “E”.

All three spend all their time washing cars, => it is on the PPF, => it’s an efficient point, => “E”.

All three spend half their time on each activity, => it shows the point “F” which is not on the PPF, => an inefficient point, => “I”.

Larry spends half his time on each activity, while Moe only washes cars and Curly only mows lawns, => it shows the point “C” on the PPF, => an efficient point.